OpenStudy (anonymous):

help with creating equations for a horizontal ellipse and a horizontal hyperbola?

OpenStudy (anonymous):

There are two fruit trees located at (3,0) and (–3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (anonymous):

I honestly did not get this lesson at all. Please explain how i can do this? Just get me to the point of having the equations written out and i should be ok. Please help :/ @jim_thompson5910 @phi

jimthompson5910 (jim_thompson5910):

They give you the focii, but they don't give you a point on the ellipse or hyperbola

OpenStudy (anonymous):

that's all i have. can it not be solved without?

jimthompson5910 (jim_thompson5910):

you'd have to know at least one point on the ellipse to find the equation of it

OpenStudy (phi):

It looks like you can choose any a, b so that it works

OpenStudy (anonymous):

ok. so i can use any point?

OpenStudy (phi):

focus is at (3,0) and (–3, 0) the center is (0,0)

jimthompson5910 (jim_thompson5910):

there are infinitely many ellipses that satisfy those 3 conditions

jimthompson5910 (jim_thompson5910):

if you made something like (5,0) a point on the ellipse, then it would uniquely determine it

OpenStudy (phi):

c is the distance from the center to one of the focus. c = 3 we use c^2 = a^2 + b^2 or 9 = a^2- b^2 where a and b are the vertex distances from the center

OpenStudy (anonymous):

so could a or b be anything that would fit in the equation acurately

jimthompson5910 (jim_thompson5910):

I think it's c^2 = a^2 - b^2

OpenStudy (phi):

yes. I would pick a=5 and b=4

jimthompson5910 (jim_thompson5910):

'a' is the length of the semi-major axis

OpenStudy (anonymous):

ok. 5 and 4 would make it easy enough.

OpenStudy (anonymous):

What about the hyperbola?

OpenStudy (phi):

see http://www.purplemath.com/modules/hyperbola.htm now we want c^2 = a^2 + b^2 c=3, so we want 9 = a^2 + b^2

OpenStudy (anonymous):

these variables represent the same things?

OpenStudy (phi):

see the link. a is the distance from the center to the vertex

OpenStudy (anonymous):

oh ok. sorry

OpenStudy (phi):

and they go into the equation x^2/a^2 - y^2/b^2 = 1

OpenStudy (anonymous):

ok. what should i do with the variables?

OpenStudy (phi):

Here is a graph of both

OpenStudy (phi):

you pick the location of the vertex (it will be between 0,0 and 3,0) i.e. the x value of the vertex will be between 0 and 3. In the plot, I picked 2 that makes a=2 and a^2 = 4 then use a^2 + b^2 = c^2 where c is 3 and a is 2 to find b, and then b^2 the equation will then be x^2/a^2 - y^2/b^2 = 1

OpenStudy (anonymous):

ok. i'm happy with those choices. Thank you so much for bringing me to these equations.